# Distilling Multi-Level X-vector Knowledge for Small Footprint Speaker Verification

### 15 Dec 2023 v2

@@ Abstract
Even though deep speaker models have demonstrated impressive accuracy in speaker verification tasks, this often comes at the expense of increased model size and computation time, presenting challenges for deployment in resource-constrained environments. Our research focuses on addressing this limitation through the development of small footprint deep speaker embedding extraction using knowledge distillation. While previous work in this domain has concentrated on speaker embedding extraction at the utterance level, our approach involves amalgamating embeddings from different levels of the x-vector model (teacher network) to train a compact student network. The results highlight the significance of frame-level information, with the student models exhibiting a remarkable size reduction of 85%-91% compared to their teacher counterparts, depending on the size of the teacher embeddings. Notably, by concatenating teacher embeddings, we achieve student networks that maintain comparable performance to the teacher while enjoying a substantial 75% reduction in model size. These findings and insights extend to other x-vector variants, underscoring the broad applicability of our approach.
Index Terms:
@@
Speaker verification, Small footprint model, Knowledge Distillation.
I Introduction
@@
Automatic speaker verification (ASV) [1] aims at identifying individuals using their voices. Advancing upon traditional, shallow statistical comparators [2, 3, 4], Gaussian mixture models (GMM) [5] and i-vectors [6], deep neural networks (DNNs) have emerged as the modern ASV solution [1]. Currently, DNNs are the predominant choice for representing speakers using speaker embeddings — fixed-dimensional vectors containing speaker-specific traits.
@@
Meanwhile, proliferating application of speech processing and biometric algorithms onto embedded devices [7, 8, 9] with constrained local computational resources has been emerged. Representative examples are smart speakers such as Amazon Alexa [^1] https://developer.amazon.com/en-US/alexa and Google Home [^2] https://home.google.com/, and assistance devices such as portable voice translator [^3] https://global.iflytek.com/jarvisen/. DNNs have various constraints to their practical deployment on them, including run-time (in)efficiency, power consumption, and memory usage due to a large number of parameters [10]. This is a major shortcoming when the embedded device has limited memory space or needs to operate with weak or no online access, ruling out cloud-based solutions. Therefore, lightweight on-device models are preferable — though this usually comes with a trade-off in recognition accuracy. Reducing the performance gap between the small and large DNN models is an important, yet challenging task in speech processing [11, 12, 13].
@@
Therefore, in this paper, we aim at reducing this gap to achieve acceptable trade-off between the model size and performance. In particular, we propose a simple and practical approach for creating a small ‘student’ DNN from a larger ‘teacher’ counterpart through knowledge distillation (KD). We review the relevant related work on small footprint ASV and KD in Sections II and III, respectively. While our focus is on ASV using de facto off-the-shelf speaker encoders, the proposed distillation framework can readily be extended to other networks and tasks (as will be discussed in Section V-B).
@@
While KD itself is not new to speech processing tasks [14, 15, 16], the practical implementation details are crucial for achieving desired model size–accuracy trade-off for a given model and task. We put forward a novel perspective on multi-level speaker representation suited for small footprint ASV. Given that the multiple layers in a DNN represent different types of features—ranging from low-level short-term spectral descriptors to abstract utterance-level latent speaker traits—a key question concerns the usefulness of the information represented by the different layers of the teacher network. While the common approach is to extract a single, utterance-level speaker embedding, we hypothesize that embeddings extracted from all the layers could help in reducing the performance gap of the teacher and the student networks. After formulating our approach in Section IV, we seek to confirm the above hypothesis experimentally in Section V. To this end, we investigate the effect of embedding-level composite (using simple concatenation) and extend the proposed methods to more advanced x-vector variants including the ones described in [17, 18, 19]. To the best of our knowledge, this is the first work that reveals the efficacy of an ‘extended’ set of multi-level speaker embeddings for developing practical, small footprint ASV systems.
@@
II Small Footprint Speaker Verification
@@
Modern neural networks for ASV are typically trained on thousands of speakers and comprise millions of learnable parameters. Despite impressive performance on common ASV evaluation corpora, the size and complexity of these models pose deployment challenges on embedded devices with limited memory and computational power. This challenge is highlighted not only for training but also, more importantly, during inference, which makes modern ASV systems be sometimes accessible only through cloud-based services with strong hardware backbones. In environments with weak or non-existent internet connection, a preference may arise for more lightweight variants of these systems, thereby alleviating the constraints on processing power and memory resources. Meanwhile, there has been a consistently-observed trade-off between model complexity and performance [20, 13]. Therefore, it becomes important to ensure that down-scaled models maintain performance levels comparable to their larger counterparts.
@@
Solutions to this problem can be divided into two broad categories. The first one, model down-scaling, includes training more compact networks [21, 22], or quantizing the model [23]. For instance, [24] proposed parameter binarization for ASV models and [25] reduced the size of self-supervised speech models via parameter sharing. While achieving promising performance in particular tasks, these approaches may demand extensive engineering effort and subtle learning schemes such as model quantization and parameter tuning. The second category, knowledge distillation (KD), also known as teacher-student (TS) learning [10][16][26], transfers the knowledge from a teacher network (usually a pre-trained DNN) to a new, student network. Compared to the first category, KD possesses task-specific design of student network and knowledge [13, 27] and tends to require less hyperparameter tuning — especially when a pre-trained teacher model is readily available (e.g. [18] [^4] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb, [19] [^5] https://github.com/yuyq96/D-TDNN). Nevertheless, producing effective knowledge for student learning remains challenging [16], especially when the student network is expected to be significantly smaller than the teacher.
@@
Having been convinced on the overall positive prospects and simplicity of the generic KD framework (stemming partly from earlier hands-on experience of the first author [28]), our work seeks to develop a novel, small footprint ASV model using KD. When DNNs were initially used for speaker embedding extraction, an open question was how the embedding should be extracted [29] — in particular, which DNN layer(s) would be most useful for speaker characterization. Models such as an x-vector with time-delay neural network (TDNN) use utterance-level (referred to as ‘segment-level’ in [30]) speaker embeddings, which contains abstract high-level latent speaker information. Meanwhile, the frame-level information at earlier layers can still be useful and its successful application as bottleneck features has been studied [31, 32], which motivates part of the methodology of this work.
@@ II-A Model Quantization
A pragmatic approach to reduce the size of a large DNN to deployable sizes involves quantizing the model itself. This option is not only practical but also reasonable, given the redundancies inherent in DNNs during training, including ineffective connections and void parameters. Recent efforts, including those focused on state-of-the-art DNN speaker encoders [18], have delved into visualizing receptive feature maps from various layers to understand their functions.
@@
One strategy to quantize the network is to prune the parameters. Noteworthy contributions in this domain, as exemplified by [33, 34], entail the removal of weights and associated connections through either elementary pruning followed by subsequent quantization and Huffman encoding, or the application of weight clustering into discrete groups represented by centroids. The latter methodology, through the transformation of centroid values, facilitates resource-efficient weight sharing while concurrently mitigating network sparsity. In the realm of convolutional neural networks (CNN), the implementation of pruning across specific dimensions within the convolutional layers has been explored [35]. Concurrently, studies have been conducted to investigate error minimization in this context [36]. A recent, promising approach involves binary (1-bit) quantization of weights [37], applied to speaker embedding extraction through conventional time-delayed neural networks (TDNNs) [38]. Alternatively, eliminating redundant parameters can be done via low-rank parameter factorization [39] achieved by estimating the importance of weight parameters through matrix decomposition techniques.
@@
Additionally, enhancing the processing efficiency of convolutional filter kernels to decrease computational complexity is another viable option. This can be achieved through factorizing the convolution kernel to two separate, cascaded kernels, with significant size reduction [40][41]. Another strategy involves splitting the filters across different channels, followed by shuffling and re-grouping, a methodology illustrated in works such as [42][43].
II-B Knowledge Distillation
@@
When implementing knowledge distillation, a certain level of information is extracted from a large ensemble of models as pseudo data distribution, which is then utilized to train a DNN model. The distribution can be generated from unlabeled data using various methods, including random sampling, mixture modeling, and designated algorithms [10].
@@
Originally applied to speech processing tasks [14][15], KD served primarily as a means to effectively transfer the knowledge of a well-trained network to a new one—but without necessarily prioritizing model size reduction. For instance, in [14], KD was employed as a domain adaptation technique, where the training data for the teacher network constituted the source domain, and the training data for the student network constituted the target domain. This approach is further explored in [15], where a semi-supervised objective based on lattice-free maximum mutual information was proposed. Both studies concentrate on speech recognition, with the training target being the state estimates of the Hidden Markov model (HMM), typically designed as the training target for the acoustic model.
III Knowledge Distillation for Speaker Verification
@@
III-A X-Vector
Given its efficacy and transparency among existing deep neural network (DNN) models [1], we consider the widely recognized vanilla x-vector speaker encoder [30, 44]. This model is founded on the time-delay neural network (TDNN) architecture [45], as depicted in Fig. 2. Selected for its efficiency in terms of model parameters, the vanilla x-vector model comprises five frame-level TDNN layers, a subsequent pooling layer, two utterance-level fully-connected layers, and an output layer. This choice allows us to explore the nuances of minimal model parameter compression using a common student network. Following the common practice, compared to the original vanilla x-vector, we substitute statistical pooling with attentive statistics pooling [46], and employ additive angular margin softmax (AAM softmax, as depicted in Fig. 2) [47] as the loss function, deviating from the conventional softmax cross-entropy. While we are aware of various data augmentation techniques [48, 49] and more advanced models [18, 19] (addressed in section VI), we choose not to factor in the combination of the influence of data, teacher model complexity, and the effectiveness of the proposed simple learning framework in the main body of this study.
III-B The Learning Framework of Knowledge Distillation
Originating in [10] and further explored in [50], knowledge distillation has recently gathered significant attention. This technique harnesses the robust representation capabilities of a well-trained, large-scale neural network to impart knowledge to a smaller counterpart. In speech and speaker recognition, the teacher network is typically trained using standard objectives like cross-entropy on a massive training corpus. Subsequently, the student network assimilates knowledge by extracting the “target distribution data” from the teacher [13, 51, 27], as illustrated in the upper part of Fig. 1. The primary objective is to align this distribution closely with the task-specific target distribution. In the case of ASV, the latter involves either speaker posterior probabilities or the direct use of speaker embedding vectors. We categorize these as posterior- and embedding-level information, respectively.
We integrate embedding-level KD, following the design outlined in [51] and [52], with slight adjustments tailored to our systems. The framework is illustrated in the lower part of Fig. 1. Informed by preliminary experiments, we opt for cosine similarity [53] as the designated loss function, defined as follows:
@@ III-B The Learning Framework of Knowledge Distillation
Figure 2: The x-vector model and the various types of embedding that are extracted from its layers. The TDNN and fully connected layers, if not marked, are all 512-dimensional. ‘AAM softmax’ denotes additive angular margin softmax [47]. The pooling layer on the left-hand side can be either statistics pooling or learnable dictionary encoding (LDE).
IV Embeddings for Knowledge Distillation
While the use of multiple levels of speaker information [54] and KD [27, 26] have been addressed in prior work, their combination for small footprint ASV demands a further investigation.
In this section, we present our methodology by examining different types of speaker embeddings for training student networks. The speaker embeddings are derived exclusively from the teacher network, as illustrated in Fig. 2. Notably, these design choices can be seamlessly transposed onto various other DNN architectures for ASV [1].
IV-A Utterance-level Embeddings
The prevalent type of speaker embedding in deep speaker verification systems is utterance-level embedding [30, 44]. For the x-vector model, we extract these embeddings, one per utterance, from the output of either the first or the second fully-connected layer following the statistics pooling layer. As indicated by [44] and confirmed by our preliminary experiments, there is no substantial performance difference between extracting from the first or the second layer. We opt for the former and term these embeddings as utterance in the experimental section.
To synchronize various levels of information, we adhere to the approach described in [13] through the following steps: 1) During the training of the student network, we replicate the speaker embedding $T$ times to generate frame-level target distribution ($T$ is set to 2 seconds for all experiments); 2) When extracting embeddings from the student network, we compute the average of the frame-level information across the input utterance.
IV-B Frame-level Embeddings
In the x-vector, the layers preceding the statistics pooling step encapsulate various frame-level information, while the subsequent layers aggregate this information into an utterance-level representation [30]. Despite the prevalence and popularity of x-vector variants, where utterance-level embeddings are the norm, the potential of frame-level information at early layers has been relatively overlooked. Motivated by earlier investigations into bottleneck (BN) features [31, 32], we explore the efficacy of frame-level embeddings as the distribution for student network learning. Specifically, we investigate bottleneck embeddings extracted from the last two TDNN layers of the teacher network, with the embeddings obtained by averaging over frame samples. The two types of embeddings are as follows:

### 19 Dec 2023 v3

# Distilling Multi-Level X-vector Knowledge for Small-footprint Speaker Verification

@@ Abstract
Even though deep speaker models have demonstrated impressive accuracy in speaker verification tasks, this often comes at the expense of increased model size and computation time, presenting challenges for deployment in resource-constrained environments. Our research focuses on addressing this limitation through the development of small-footprint deep speaker embedding extraction using knowledge distillation. While previous work in this domain has concentrated on speaker embedding extraction at the utterance level, our approach involves amalgamating embeddings from different levels of the x-vector model (teacher network) to train a compact student network. The results highlight the significance of frame-level information, with the student models exhibiting a remarkable size reduction of 85%-91% compared to their teacher counterparts, depending on the size of the teacher embeddings. Notably, by concatenating teacher embeddings, we achieve student networks that maintain comparable performance to the teacher while enjoying a substantial 75% reduction in model size. These findings and insights extend to other x-vector variants, underscoring the broad applicability of our approach.
keywords:
@@
Speaker recognition, Small foot-print model, Knowledge distillation
journal: Nuclear Physics B
1 Introduction
@@
Automatic speaker verification (ASV) dnn_asv2021 aims at identifying individuals using their voices. Advancing upon traditional, shallow statistical comparators simiarity_asv2010 , Gaussian mixture models (GMM) gmm ; gmm_earlyasv_2008 and i-vectors ivector ; ivector_asv2019 , deep neural networks (DNNs) have emerged as the modern ASV solution dnn_asv2021 . Currently, DNNs are the predominant choice for representing speakers using speaker embeddings — fixed-dimensional vectors containing speaker-specific traits.
@@
Meanwhile, proliferating applications of speech processing and biometric algorithms onto embedded devices assistant1 ; assistant2 ; assistant3 with constrained local computational resources have emerged. Representative examples are smart home speakers such as Amazon Alexa [^1] https://developer.amazon.com/en-US/alexa and Google Home [^2] https://home.google.com/, and assistance devices such as portable voice translator [^3] https://global.iflytek.com/jarvisen/. DNNs have various constraints to their practical deployment on them, including run-time (in)efficiency, power consumption, and memory usage due to a large number of parameters kd2006 . This is a major shortcoming when the embedded device has limited memory space or needs to operate with weak or no online access, ruling out cloud-based solutions. Therefore, lightweight on-device models are preferable — though this usually comes with a trade-off in recognition accuracy. Reducing the performance gap between the small and large DNN models is an important, yet challenging task in speech processing ondevice3 ; ondevice1 ; tslearning_slt2018 .
@@
Therefore, in this paper, we aim at reducing this gap to achieve acceptable trade-off between the model size and performance. In particular, we propose a simple and practical approach for creating a small ‘student’ DNN from a larger ‘teacher’ counterpart through knowledge distillation (KD). We review the relevant related work on small footprint ASV and KD in Sections 2 and 3, respectively. While our focus is on ASV using de facto off-the-shelf speaker encoders, the proposed distillation framework can readily be extended to other networks and tasks (as will be discussed in Section 5.2).
@@
While KD itself is not new to speech processing tasks kd_asr2013 ; kd_asr2018 ; kd2021 , the practical implementation details are crucial for achieving desired model size–accuracy trade-off for a given model and task. We put forward a novel perspective on multi-level speaker representation suited for small footprint ASV. Given that the multiple layers in a DNN represent different types of features—ranging from low-level short-term spectral descriptors to abstract utterance-level latent speaker traits—a key question concerns the usefulness of the information represented by the different layers of the teacher network. While the common approach is to extract a single, utterance-level speaker embedding, we hypothesize that embeddings extracted from all the layers could help in reducing the performance gap of the teacher and the student networks. After formulating our approach in Section 4, we seek to confirm the above hypothesis experimentally in Section 5. To this end, we investigate the effect of embedding-level composite (using simple concatenation) and extend the proposed methods to more advanced x-vector variants including the ones described in Snyder_etdnn_2019 ; ecapa ; dtdnn . To the best of our knowledge, this is the first work that reveals the efficacy of an ‘extended’ set of multi-level speaker embeddings for developing practical, small footprint ASV systems.
@@
2 Small Footprint Speaker Verification
@@
Modern neural networks for ASV are typically trained on thousands of speakers and comprise millions of learnable parameters. Despite impressive performance on common ASV evaluation corpora, the size and complexity of these models pose deployment challenges on embedded devices with limited memory and computational power. This challenge is highlighted not only for training but also, more importantly, during inference, which makes modern ASV systems be sometimes accessible only through cloud-based services with strong hardware backbones. In environments with weak or non-existent internet connection, a preference may arise for more lightweight variants of these systems, thereby alleviating the constraints on processing power and memory resources. Meanwhile, there has been a consistently-observed trade-off between model complexity and performance energy_efficient_net2018 ; tslearning_slt2018 . Therefore, it becomes important to ensure that down-scaled models maintain performance levels comparable to their larger counterparts.
@@
Solutions to this problem can be divided into two broad categories. The first one, model down-scaling, includes training more compact networks mobilenet ; efficientnet , or quantizing the model quantized_net . For instance, binarized_net_asv2021 proposed parameter binarization for ASV models and audio_albert reduced the size of self-supervised speech models via parameter sharing. While achieving promising performance in particular tasks, these approaches may demand extensive engineering effort and subtle learning schemes such as model quantization and parameter tuning. The second category, knowledge distillation (KD), also known as teacher-student (TS) learning kd2006 ; kd2021 ; kd2022 , transfers the knowledge from a teacher network (usually a pre-trained DNN) to a new, student network. Compared to the first category, KD possesses task-specific design of student network and knowledge tslearning_slt2018 ; kd2020 and tends to require less hyperparameter tuning — especially when a pre-trained teacher model is readily available (e.g. ecapa  [^4] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb, dtdnn  [^5] https://github.com/yuyq96/D-TDNN). Nevertheless, producing effective knowledge for student learning remains challenging kd2021 , especially when the student network is expected to be significantly smaller than the teacher.
@@
Having been convinced on the overall positive prospects and simplicity of the generic KD framework (stemming partly from earlier hands-on experience of the first author kd_asv2018 ), our work seeks to develop a novel, small footprint ASV model using KD. When DNNs were initially used for speaker embedding extraction, an open question was how the embedding should be extracted xvector_2016 — in particular, which DNN layer(s) would be most useful for speaker characterization. Models such as an x-vector with time-delay neural network (TDNN) use utterance-level (referred to as ‘segment-level’ in xvector2017 ) speaker embeddings, which contains abstract high-level latent speaker information. Meanwhile, the frame-level information at earlier layers can still be useful and its successful application as bottleneck features has been studied dvector ; bnf , which motivates part of the methodology of this work.
@@ 2.1 Model Quantization
A pragmatic approach to reduce the size of a large DNN to deployable sizes involves quantizing the model itself. This option is not only practical but also reasonable, given the redundancies inherent in DNNs during training, including ineffective connections and void parameters. Recent efforts, including those focused on state-of-the-art DNN speaker encoders ecapa , have delved into visualizing receptive feature maps from various layers to understand their functions.
@@
One strategy to quantize the network is to prune the parameters. Noteworthy contributions in this domain, as exemplified by deep_compression ; deep_compression2015 , entail the removal of weights and associated connections through either elementary pruning followed by subsequent quantization and Huffman encoding, or the application of weight clustering into discrete groups represented by centroids. The latter methodology, through the transformation of centroid values, facilitates resource-efficient weight sharing while concurrently mitigating network sparsity. In the realm of convolutional neural networks (CNN), the implementation of pruning across specific dimensions within the convolutional layers has been explored cnn_compression2017 . Concurrently, studies have been conducted to investigate error minimization in this context cnn_compression2016 . A recent, promising approach involves binary (1-bit) quantization of weights binarized_nnet2021 , applied to speaker embedding extraction through conventional time-delayed neural networks (TDNNs) binarized_nnet2020 . Alternatively, eliminating redundant parameters can be done via low-rank parameter factorization cnn_decomposition2015 achieved by estimating the importance of weight parameters through matrix decomposition techniques.
@@
Additionally, enhancing the processing efficiency of convolutional filter kernels to decrease computational complexity is another viable option. This can be achieved through factorizing the convolution kernel to two separate, cascaded kernels, with significant size reduction mobilenet_v2 ; mobilenet_asv2020 . Another strategy involves splitting the filters across different channels, followed by shuffling and re-grouping, a methodology illustrated in works such as shufflenet ; shufflenet_v2 .
2.2 Knowledge Distillation
@@
When implementing knowledge distillation, a certain level of information is extracted from a large ensemble of models as pseudo data distribution, which is then utilized to train a DNN model. The distribution can be generated from unlabeled data using various methods, including random sampling, mixture modeling, and designated algorithms kd2006 .
@@
Originally applied to speech processing tasks kd_asr2013 ; kd_asr2018 , KD served primarily as a means to effectively transfer the knowledge of a well-trained network to a new one—but without necessarily prioritizing model size reduction. For instance, in kd_asr2013 , KD was employed as a domain adaptation technique, where the training data for the teacher network constituted the source domain, and the training data for the student network constituted the target domain. This approach is further explored in kd_asr2018 , where a semi-supervised objective based on lattice-free maximum mutual information was proposed. Both studies concentrate on speech recognition, with the training target being the state estimates of the Hidden Markov model (HMM), typically designed as the training target for the acoustic model.
3 Knowledge Distillation for Speaker Verification
@@
3.1 X-Vector
Given its efficacy and transparency among existing deep neural network (DNN) models dnn_asv2021 , we consider the widely recognized vanilla x-vector speaker encoder xvector2017 ; xvector2018 . This model is founded on the time-delay neural network (TDNN) architecture tdnn2015 , as depicted in Fig. 2. Selected for its efficiency in terms of model parameters, the vanilla x-vector model comprises five frame-level TDNN layers, a subsequent pooling layer, two utterance-level fully-connected layers, and an output layer. This choice allows us to explore the nuances of minimal model parameter compression using a common student network. Following the common practice, compared to the original vanilla x-vector, we substitute statistical pooling with attentive statistics pooling astats_pooling , and employ additive angular margin softmax (AAM softmax, as depicted in Fig. 2) aam_softmax as the loss function, deviating from the conventional softmax cross-entropy. While we are aware of various data augmentation techniques rir ; musan and more advanced models ecapa ; dtdnn (addressed in section 6), we choose not to factor in the combination of the influence of data, teacher model complexity, and the effectiveness of the proposed simple learning framework in the main body of this study.
3.2 The Learning Framework of Knowledge Distillation
Originating in kd2006 and further explored in kd2015 , knowledge distillation has recently gathered significant attention. This technique harnesses the robust representation capabilities of a well-trained, large-scale neural network to impart knowledge to a smaller counterpart. In speech and speaker recognition, the teacher network is typically trained using standard objectives like cross-entropy on a massive training corpus. Subsequently, the student network assimilates knowledge by extracting the “target distribution data” from the teacher tslearning_slt2018 ; ts_asr_slt2018 ; kd2020 , as illustrated in the upper part of Fig. 1. The primary objective is to align this distribution closely with the task-specific target distribution. In the case of ASV, the latter involves either speaker posterior probabilities or the direct use of speaker embedding vectors. We categorize these as posterior- and embedding-level information, respectively.
We integrate embedding-level KD, following the design outlined in ts_asr_slt2018 and tslearning2019 , with slight adjustments tailored to our systems. The framework is illustrated in the lower part of Fig. 1. Informed by preliminary experiments, we opt for cosine similarity cosine_similarity as the designated loss function, defined as follows:
@@ 3.2 The Learning Framework of Knowledge Distillation
Figure 2: The x-vector model and the various types of embedding that are extracted from its layers. The TDNN and fully connected layers, if not marked, are all 512-dimensional. ‘AAM softmax’ denotes additive angular margin softmax aam_softmax . The pooling layer on the left-hand side can be either statistics pooling or learnable dictionary encoding (LDE).
4 Embeddings for Knowledge Distillation
While the use of multiple levels of speaker information aggregated_stats_2020 and KD kd2020 ; kd2022 have been addressed in prior work, their combination for small footprint ASV demands a further investigation.
In this section, we present our methodology by examining different types of speaker embeddings for training student networks. The speaker embeddings are derived exclusively from the teacher network, as illustrated in Fig. 2. Notably, these design choices can be seamlessly transposed onto various other DNN architectures for ASV dnn_asv2021 .
4.1 Utterance-level Embeddings
The prevalent type of speaker embedding in deep speaker verification systems is utterance-level embedding xvector2017 ; xvector2018 . For the x-vector model, we extract these embeddings, one per utterance, from the output of either the first or the second fully-connected layer following the statistics pooling layer. As indicated by xvector2018 and confirmed by our preliminary experiments, there is no substantial performance difference between extracting from the first or the second layer. We opt for the former and term these embeddings as utterance in the experimental section.
To synchronize various levels of information, we adhere to the approach described in tslearning_slt2018 through the following steps: 1) During the training of the student network, we replicate the speaker embedding $T$ times to generate frame-level target distribution ($T$ is set to 2 seconds for all experiments); 2) When extracting embeddings from the student network, we compute the average of the frame-level information across the input utterance.
4.2 Frame-level Embeddings
In the x-vector, the layers preceding the statistics pooling step encapsulate various frame-level information, while the subsequent layers aggregate this information into an utterance-level representation xvector2017 . Despite the prevalence and popularity of x-vector variants, where utterance-level embeddings are the norm, the potential of frame-level information at early layers has been relatively overlooked. Motivated by earlier investigations into bottleneck (BN) features dvector ; bnf , we explore the efficacy of frame-level embeddings as the distribution for student network learning. Specifically, we investigate bottleneck embeddings extracted from the last two TDNN layers of the teacher network, with the embeddings obtained by averaging over frame samples. The two types of embeddings are as follows:
  • 1.